fraction 4/x = 5/7 solve and check work

Answer by checkley77(12844)

4/x = 5/7 CROSS MULTIPLY
5X=4*7
5X=28
X=28/5 ANSWER.
PROOF:
4/(28/5)=5/7
5(28/5)=4*7
28=28